# Should AI Optimize Your Code? A Comparative Study of Current Large Language Models Versus Classical Optimizing Compilers

### v1

@@ 4. Evaluation: LLMs versus Classical Optimizing Compilers
This section assesses the current capabilities of two Large Language Models, GPT-4.0 and CodeLlama-70B, in comparison with three different automatic compilers, CETUS, PLUTO and ROSE. Section  4.1 outlines the experimental setup, Section  4.2 evaluates the ability of the LLMs to process and correctly optimize large code sizes. Section  4.3 examines the ability of these LLMs to optimize code sections effectively and also assesses their capacity in handling the challenge patterns introduced in Section 3.2. Lastly, Section  4.4 compares the achieved speedup of all five Automatic Optimization Tools (AOTs).
@@ 4.1. Experimental setup
To assess our Automatic Optimization Tools, we use the PCAOT environment described in Section  3.1. Our test suite drew from three distinct benchmarks: The NPB v3.3 (Bailey et al., 1991) , the PB suite V4.2  (Yuki and Pouchet, [n. d.]), and the PCB benchmark introduced in Section 3.2. For our experiments, we measured the performance of the applications using input Class B for the NPB and the LARGE_DATASET for the PB and PCB suite.
We have selected specific subroutines from the NPB and the PB suites that represent challenge patterns. For the PCB, we developed 20 use cases, with each one representing one of the six challenge scenarios detailed in Section 3.2.
Table  1 displays our test suite. We set the following AOT parameters: For the CETUS, PLUTO and ROSE compilers we chose their default options, selecting the known, safe optimization behavior. For the LLMs, we chose a Temperature of 0.2 and a Top_p value of 0.1, increasing deterministic behavior for the code generation tasks.
Table 1. Testing Dataset: Experimental Section, Impact of the experimental section on the overall application, Pattern evaluated, and number of loops we are testing with the AOTs
Benchmark Application Name Experimental Section Impact % Pattern evaluated #Loops
NAS BT initialize 0.2% PF, PR 11
NAS BT rhs_norm 0.5% PR, PO, NW 1
NAS BT compute_rhs 10.2% PR, PO 11
NAS BT x_solve 27.7% PF, PO 1
NAS SP initialize 0.3% PF, PR 11
NAS SP rhs_norm 0.2% PF, PR 1
NAS SP y_solve 19.6% PF, PO 1
NAS SP exact_rhs 0.3% PR, PO, NW 5
NAS MG norm2u3 1.1% PF, PR, NW 1
NAS MG zran3 29.3% PF, PR 3
NAS CG conj_grad 98.9% PR, PA, PO, NW 5
NAS IS full_verify 99.8% DS, PO 5
Poly Linear Algebra 2mm 91.8% PR, PO 1
Poly Linear Algebra 3mm 98.4% PR, PO 3
Poly Linear Algebra Atax 18.2% PR, PO 5
Poly Datamining Correlation 99% PR, PO 1
PCB Challenges PO_Function_Version[1-5] 97% PO 5
PCB Challenges PF_Function_Version[1-3] 93% PF 4
PCB Challenges PR_Function_Version[1-3] 98% PR 9
PCB Challenges DS_Function_Version[1-3] 96.3% DS 3
PCB Challenges PA_Function_Version[1-3] 95% PA 4
PCB Challenges NW_Function_Version[1-3] 97.2% NW 3
We measured execution times on a compute node featuring a 4-core Intel Xeon Gold 6230 processor configuration in dual sockets. Each processor operates at a base frequency of 2.1 GHz, with a 27.5MB cache, and was supported by up to 1 GB of DDR4 memory. Application codes were compiled using GCC v4.8.5 with -O3 optimization on CentOS v7.4.1708. Reported values represent the median of three application runs, each utilizing one thread per core.
@@ 4.2. Impact of Code Size
Table  2, shows the capacity of the Large Language Models to process programs with large code sizes. This table displays the number of lines of code per experimental section versus the successful cases after sending the two different prompts described in Section  3.1.1, 3 times each to the two chosen LLMs; GTP4.0 and Codellama-70B, with a total of 192 attempts.
These results illustrate the capability of the LLMs in handling large code sizes. While smaller code sections tend to be transformed more accurately, they are not guaranteed to be 100% correct. As the code size increases, the likelihood of inaccuracies also increases.
Program sections of approximately 20 lines often produce correct output, whereas no section above 200 lines was transformed successfully (performance will be evaluated later). In between, the LLMs exhibit non-deterministic behavior, with some runs producing useful results while others with the same input fail. GPT4.0 and Codellama-70B return successfully transformed code in 14.06% of the cases.
These results further justify our method of employing LLM capabilities on the basis of individual program sections.
@@ 4.3. LLM Optimization Capabilities
Figure 9 shows the success rates of generating correctly optimized code by the tested LLMs on specific challenge patterns, using the two prompts described in Section 3.1.1. The results are classified into four distinct categories: Pattern correctly applied, Pattern not correctly applied but code compiles, Compilation error, Incorrect Result and Runtime error.
Table 2. Capacity of Large Language Models to process large programs
Large Language Model
GPT4.0 CodeLlama-70B
Benchmark
Experimental
Section
#Lines
code
Succesfull
Cases
Benchmark
Experimental
Section
#Lines
code
Sucessfull
cases
BT
Linear Algebra
Datamining
SP
MG
Linear Algebra
Linear Algebra
IS
BT
SP
CG
MG
SP
SP
BT
BT
rhs_norm
Atax
Correlation
rhs_norm
norm2u3
2mm
3mm
full_verify
Initialize
Initialize
conj_grand
zran3
y_solve
exact_rhs
x_solve
compute_rhs
18
21
21
26
28
29
34
41
155
161
183
187
250
313
346
382
3
5
5
0
2
6
6
0
0
0
0
0
0
0
0
0
BT
Linear Algebra
Datamining
SP
MG
Linear Algebra
Linear Algebra
IS
BT
SP
CG
MG
SP
SP
BT
BT
rhs_norm
Atax
Correlation
rhs_norm
norm2u3
2mm
3mm
full_verify
Initialize
Initialize
conj_grand
zran3
y_solve
exact_rhs
x_solve
compute_rhs
18
21
21
26
28
29
34
41
155
161
183
187
250
313
346
382
2
5
5
0
0
6
6
0
0
0
3
0
0
0
0
0
Figure 9. Correctness Evaluation for the following challenging patterns: Array Reduction, NOWAIT, Parallel-Regions, Parallelization at the Outermost level (P-Outermost), Parallelization for function calls (P-Function Calls), Dynamic Scheduling (Dynamic Sch)
GPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IPGPT4-COTGPT4-IPCLLama-70B-CoTCLLama-70B-IPGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70-IPGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70-IPGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IPGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IP002020202040404040606060608080808010010010010025252525% $41.7$% $41.7$% $15.2$% $31.8$% $28.8$% $21.7$% $20.2$% $25$% $25$% $15.2$% $21.2$% $18.6$% $20.9$% $31.8$% $20.2$% $18.7$% $17.3$% $25$% $72.7$% $71.2$% $47$% $59.1$% $63.6$% $64.3$% $44.2$% $57.4$% $33.3$% $75$% $33.3$% $75$% Array ReductionNOWAITParallel-RegionsP-OutermostP-Fuction CallsDynamic SchPercentagePattern Correctly Applied (%)Pattern No Correctly Applied but Compiled(%)Compilation Error (%)Incorrect Code(%)Runtime Error (%)
Our primary finding from this information is that LLMs are not yet in a position where they can serve as automatic optimizers. Only in three of the analyzed optimization patterns, both LLMs produce partly incorrect results. A key demand on automatic optimizers is that they be 100% correct; tools that require software engineers to engage in possibly lengthy debug phases after use are not acceptable. The partial successes in two of the patterns (Array Reduction and Parallelizing Outermost Loop) indicate that LLMs may be making progress toward the goal of becoming useful AOTs. These two patterns represent some of the most important parallelization capabilities and likely have many code examples serving as LLM training data, which may explain the partial success. The NOWAIT and Parallel-Region patterns tend to involve large code regions, which challenge LLMs, as discussed in Section 4.2. The two LLMs completely fail on Dynamic Scheduling patterns. We attribute this behavior to the more complex code exhibited by such test cases. Dynamic scheduling tends to be needed in loops that involve irregular code and data structures, which are more difficult to comprehend.
The results show no significant difference between the two prompting methods we used. While other researchers have reported more success with Chain-of-Thought (CoT) prompting, this method performed marginally worse in some cases than the more direct Instruction Prompting (IP) method. Finally, comparing the two LLMs shows insignificant differences as well. CodeLLama is optimized for program code synthesis, as well as infilling/completion  (Grattafiori et al., 2023), hence we expected it to perform well. However, ChatGPT has a larger number of AI model parameters, which may make up for the fact that it is a general model.
4.4. Speedup Evaluation
Figure 10 shows the maximum average speedups for correct and incorrect code obtained from the five distinct AOTs across our three benchmark suites against each benchmark’s baseline (Serial code) and its Hand-optimized version.
Figure 10. AOTs - Speedup Evaluation of each AOT using a specific configuration: Chain-of-Thought and Instruction prompting for LLMs and the default configuration for the optimizing compilers.
GPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IPCETUSROSEHANDGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IPCETUSPLUTOROSEGPT4-CoTGPT4-IPCLLama-70B-CoTCLLama-70B-IPCETUSROSEHAND002222444466668888x1.11.11.11.1 x1.11.11.11.1 x0.80.80.80.8 x1.31.31.31.3 x1.61.61.61.6 x1.91.91.91.9 x1111 x2.32.32.32.3 x1.61.61.61.6 x2222 x2.12.12.12.1 x2222 x1.61.61.61.6 x7777 x1.31.31.31.3 x1.41.41.41.4 x1.21.21.21.2 x1.91.91.91.9 x1.61.61.61.6 x1.51.51.51.5 x1.21.21.21.2 x2.22.22.22.2 NASPolyBenchPCBSpeed upAverage speedUpAdditional speedUp but Incorrect Code
Our findings demonstrate the efficacy of CodeLlama-70B as the superior optimizer among Language Model-based (LLM) solutions, capable of achieving speedups of up to 2.1x compared to the original program, measuring correct code only. When considering incorrectly optimized code as well, the achieved speedup is approximately 2.7x.
Conversely, we can observe a noticeable difference between all the AOTs and PLUTO in terms of speedup achieved within the PolyBenchmark suite, where PLUTO outperformed all of the AOTs with a remarkable 7x speedup. This superior performance is attributed to PLUTO’s specialization on the Polyhedral model (Bondhugula et al., 2008b). The specialization may also explain why PLUTO was unable to process the other benchmark suites.
Overall, CETUS emerged as the best among the optimizing compilers for processing a variety of programs within these suites, achieving a maximum speedup of 1.9x. The ROSE optimizer achieved a speedup across all benchmarks of 1.2x, remaining below the other AOTs in terms of efficacy.
While GPT4 has demonstrated success across various tasks, it exhibited the lowest performance for code optimization across all the benchmarks and among the LLMs. It achieved a maximum speedup of 1.6x, falling short of CodeLlama-70B.
@@ 5. Related work
The primary objective of PCAOT is to automatically validate the correctness and performance of optimizations generated by Automatic Optimization Tools (AOTs). To the best of our knowledge, no other research has directly addressed this specific objective. However, several studies have created benchmarks to evaluate the capabilities of Large Language Models (LLMs) in optimization tasks.
For instance, HumanEval (Black et al., 2022) is a benchmark designed to assess the performance of LLMs in generating correct and efficient code. Additionally, there has been a study evaluating the ability of LLMs to synthesize short programs from natural language descriptions, demonstrating the potential of these models in code generation tasks (Austin et al., 2021).
Recent studies have also evaluated the performance of LLMs in specific High-Performance Computing (HPC) optimization tasks; such as OpenMP pragma prediction (Nichols et al., 2023) highlighting the application of LLMs in parallel computing optimizations. Additionally, approaches to code completion (Valero-Lara et al., 2023) and issue fixing (Valero-Lara et al., 2023) show the breadth of LLM applications in software development and maintenance.
Another significant contribution in this domain is the novel paradigm introduced by researchers using Large Language Models with compiler feedback to optimize the code size of LLVM assembly. Their model takes unoptimized LLVM Intermediate Representation (IR) as input and produces optimized IR, optimization passes, and instruction counts for both unoptimized and optimized IRs  (Grubisic et al., 2024).
Despite these advancements, the automatic validation of both the correctness and performance of optimizations remains underexplored. Our work aims to fill this gap by providing a comprehensive framework for evaluating and validating the outputs of AOTs, ensuring that the generated optimizations meet the desired standards of correctness. We also proposed a novel Parallel Computing Challenge Benchmark suite (PCB) V1.0 containing six challenge patterns for evaluating AOT code optimization capabilities.
@@ 6. Conclusions
This paper introduced a novel mechanism called PCAOT, which enables the automatic validation of the correctness and performance of Automatic Optimization Tools-generated code. While we have applied the mechanism to verify LLM-generated code, it also enables unsafe compiler optimization to be tested. Additionally, we have presented the Parallel Computing Challenge Benchmark suite (PCB) v1.0, comprising 20 use cases for evaluating the capabilities of Automatic Optimization Tools. To the best of our knowledge, this is the first such benchmark.
Our results show that LLMs only succeed in optimizing very small programs of approximately 20 lines. Given that real-world software often contains thousands of lines of code, these models are not yet ready to serve as automatic optimizers.
Our results also show that LLMs perform only marginally better than current optimizing compilers, in terms of correctness. The chosen LLMs can correctly identify three out of the six different challenge patterns, such as Array Reduction, Parallelizing Outermost Loops, and Loops Containing Function Calls, with a success rate higher than 50% in most cases. However, they often fail on other patterns, including eliminating barrier synchronization using the NOWAIT clause, parallel regions enclosing multiple parallel loops, and avoiding load imbalance through dynamic scheduling.
In terms of speedup, the best compiler and the best LLM performed within 10%. We also found no significant difference between the two prompting methods we employed, with Chain-of-Thought (CoT) performing slightly worse in some cases than the more straightforward Instruction Prompting (IP) method.
Finally, addressing the main question asked in the introduction: Can AI-driven models revolutionize the way we approach code optimization? We find that, with their current capabilities, the LLMs cannot yet suitably serve as Automatic Optimization Tools. Even though the potential exists, future LLMs will need significant improvements in terms of performance and correctness.
7. ACKNOWLEDGEMENTS

### v2

@@ 4. Evaluation: LLMs versus Classical Optimizing Compilers
This section assesses the current capabilities of two Large Language Models, GPT-4.0 and CodeLlama-70B, in comparison with three different automatic compilers, CETUS, PLUTO and ROSE. Section  4.1 outlines the experimental setup, Section  4.2 evaluates the ability of the LLMs to process and correctly optimize code of different sizes. Section  4.3 examines the ability of these LLMs to optimize code sections effectively and also assesses their capacity in handling the challenge patterns introduced in Section 3.2. Lastly, Section  4.4 compares the achieved speedup of all five Automatic Optimization Tools (AOTs).
@@ 4.1. Experimental setup
To assess our Automatic Optimization Tools, we use the PCAOT environment described in Section  3.1. Our test suite drew from three distinct benchmarks: The NPB v3.3 (Bailey et al., 1991) , the PB suite V4.2  (Yuki and Pouchet, [n. d.]), and the PCB benchmark introduced in Section 3.2. We measured the performance of the applications using input Class B for the NPB and the LARGE_DATASET for the PB and PCB suite. Recall that we are focusing on challenge patterns for the optimizing compilers. We have selected specific subroutines from the NPB and the PB suites that contains such cases with a total of 16 experimental sections. For the PCB, we developed 28 use cases, with each one representing one of the six challenge scenarios described in Section 3.2. In total, our testing dataset contains 44 experimental sections. More details about the testing data set can be found on our GitHub repository (Miguel Rosas, [n. d.]).
We set the following AOT parameters: For the CETUS, PLUTO and ROSE compilers we chose their default options, selecting the known, safe optimization behavior. For the LLMs, we chose a Temperature of 0.2 and a Top_p value of 0.1, increasing deterministic behavior for the code generation tasks.
We measured execution times using 4 cores on a compute node featuring an Intel Xeon Gold 6230 processor configuration in dual sockets. Each processor operates at a base frequency of 2.1 GHz, with a 27.5MB cache, and was supported by up to 1 GB of DDR4 memory. Application codes were compiled using GCC v11.2 with -O3 optimization on CentOS v7.4.1708. Reported values represent the median of three application runs, each utilizing one thread per core.
@@ 4.2. Impact of Code Size
Figure 3, shows the failure rate of the Large Language Models in generating programs with increasing code sizes. An unsuccessful case refers to instances where the LLMs failed to produce output that passed validation. The graph displays the number of lines of code per experimental section versus the unsuccessful cases. We sent the sections to GPT4.0 and Codellama-70B, using the three different prompts described in Section  3.1.1, repeating each run three times, with a total of 792 attempts (44 experimental sections x 3 times each x 3 prompts x 2 LLMs = 792 attempts). Details about the experimental sections used in this study can be found on our GitHub repository (Miguel Rosas, [n. d.]).
These results picture the LLMs’ capability to handle large code sizes. While smaller code sections tend to be transformed more accurately, they are not guaranteed to be 100% correct. As the code size increases, the likelihood of inaccuracies also increases.
Program sections of approximately 20 lines often produce correct output, whereas sections above 20 lines increase the failing percentage (performance will be evaluated later). In between, the LLMs exhibit non-deterministic behavior, with some runs producing useful results while others with the same input fail. GPT4.0 and Codellama-70B return successfully transformed code in 16.75% of the cases. These results corroborate our method of employing LLM capabilities on the basis of individual program sections, such as loops.
Figure 3. Average failure rate of the Large Language Models across benchmarks when generating programs with increasing code sizes.
@@ 4.3. LLM Optimization Capabilities
In this subsection, we evaluate the capacity of Large Language Models (LLMs) to optimize both parallel and non-parallel scenarios. Parallel use cases represent scenarios where the LLMs must identify opportunities for parallelization and apply the correct optimization strategy. In contrast, non-parallel use cases test the LLMs capability to accurately recognize scenarios where parallelization is not feasible or beneficial. For this purpose, we utilized 94 parallel use cases and 10 non-parallel use cases.
4.3.1. Parallel Use cases:
@@
Figure 4 shows the success rates of generating correctly optimized code by the tested LLMs on specific challenge patterns, using the three prompts described in Section 3.1.1. The results are classified into three distinct categories: Expected Pattern applied, Unexpected Pattern Correctly Applied and Errors (Compilation and Runtime Errors).
Figure 4. Correctness Evaluation for the challenge patterns: Parallelization at the Outermost level (P-Outermost), Parallelization for function calls (P-Function Calls), Array Reduction, Parallel-Regions, NOWAIT, and Dynamic Scheduling (Dynamic Sch).
Our primary finding from this information is that LLMs are not yet in a position where they can serve as automatic optimizers. Only in three of the analyzed optimization patterns, both LLMs produce partly correct results. A key demand on automatic optimizers is that they be 100% correct; tools that require software engineers to engage in possibly lengthy debug phases after use are not acceptable. The partial successes in three of the patterns (Parallelizing Outermost Loop, Parallelization with Function Calls and Array Reduction) indicate that LLMs may be making progress toward the goal of becoming useful AOTs. These three patterns represent some of the most important parallelization capabilities and likely have many code examples serving as LLM training data, which may explain the partial success. The Parallel-Region and NOWAIT patterns tend to involve large code regions, which challenge LLMs, as discussed in Section 4.2. The two LLMs completely fail on Dynamic Scheduling patterns. We attribute this behavior to the more complex code exhibited by such test cases. Dynamic scheduling tends to be needed in loops that involve irregular code and data structures, which are more difficult to comprehend.
The results show a substantial difference between the three prompting strategies, with DIP being better among the three different approaches in most of the cases. While other researchers have reported more success with CoT prompting, this method performed essentially similarly to DIP.
4.3.2. Non-parallel Use cases
Figure 5 illustrates the correctness of the non-parallel use cases. For this analysis, we evaluate loops with data dependencies (DP) and loops with non-side effect free function calls (FC).
  • Non_Parallel_Loop_DP: This category represents code sections with data dependencies that prevent safe parallelization.
  • Non_Parallel_Loop_FC: This category represents code sections containing function calls that are not side-effect-free, thereby making data dependence analysis required for safe parallelization more complex.
These categories help differentiate the specific reasons why loops in the negative use cases cannot be parallelized, assessing the LLMs’ ability to correctly recognize non-parallelizable scenarios. The results show that the DIP strategy performed marginally better than the CoT approach in recognizing non-parallel use cases, a similar result was found for the parallel use cases show in Section  4.3.1. However, both LLMs failed entirely to identify non-parallel code when using IP. We attribute this to the lack of specificity in the IP prompt, as explained in Section  3.1.1.
Figure 5. Correctness Evaluation for the challenge patterns: Parallelization at the Outermost level and Parallelization for function calls in non-parallel Scenarios. The expected behavior in this case is that the LLM recognizes the code as non parallelizable.
4.4. Speedup Evaluation by Pattern
Figure 6 shows the maximum average speedups of the correctly optimized codes for the parallel scenarios from the two chosen LLMs. These scenarios include GPT4.0 and CodeLlama-70B across our testing Dataset described in (Miguel Rosas, [n. d.]) with the three different prompts described in Section 3.1.1 for three specific patterns: Parallelization at the Outermost level, Array Reduction and Parallelization for function calls.
Figure 6. Speedup Evaluation two LLMs (AOTs) across three challenging patterns
Our results demonstrate the efficacy of CodeLlama-70B, which achieves speedups of up to 1.95x and 1.79x over the original program. Additionally, the DIP prompt is more efficient for this particular AOT. We attribute this behavior to the type of model that is being used. CodeLlama-70B is fine-tuned to follow instructions more closely. (META, [n. d.]).
4.5. Speedup Evaluation for each AOT
Figure 7 shows the maximum average speed-up of the correctly Parallel optimized codes obtained from the five distinct AOTs over the serial and the Hand-optimized versions.
Our results show CodeLlama-70B as the superior optimizer among Language Model-based (LLM) solutions, capable of achieving speedups of up to 1.75x compared to the original program.
Conversely, we can observe a noticeable difference between all AOTs and PLUTO in terms of the speedup achieved within the PB suite, where PLUTO outperformed all AOTs with a remarkable 7x speedup. This superior performance is attributed to PLUTO’s specialization on the Polyhedral model (Bondhugula et al., 2008b). The specialization may also explain why PLUTO was unable to process the other benchmark suites.
Overall, CETUS emerged as the best among the optimizing compilers for processing a variety of programs within these suites, achieving a maximum speedup of 1.67x. The ROSE optimizer achieved a speedup across all benchmarks of 1.17x, remaining below the other AOTs in terms of efficacy.
While GPT4 has demonstrated success across various tasks, it exhibited the lowest performance for code optimization across all the benchmarks and among the LLMs. It achieved a maximum speedup of 1.34x, falling short of CodeLlama-70B.
Figure 7. AOTs - Speedup Evaluation of each AOT using a specific configuration: Simple Instruction, Instruction prompting, and Chain-of-Thought for LLMs and the default configuration for the optimizing compilers.
@@ 5. Related work
Recent studies have evaluated the performance of LLMs in specific High-Performance Computing (HPC) optimization tasks, such as OpenMP pragma prediction (Nichols et al., 2023) highlighting the application of LLMs in parallel computing optimizations. Additionally, approaches to code completion (Valero-Lara et al., 2023) and issue fixing (Valero-Lara et al., 2023) show the breadth of LLM applications in software development and maintenance.
Another significant contribution in this domain is the novel paradigm introduced by researchers using LLMs with compiler feedback to optimize the code size of LLVM assembly. Their model takes unoptimized LLVM Intermediate Representation (IR) as input and produces optimized IR, optimization passes, and instruction counts for both unoptimized and optimized IRs  (Grubisic et al., 2024).
Recent studies have evaluated the ability of LLMs to generate short programs from natural language descriptions, demonstrating the potential of these models in code generation tasks (Austin et al., 2021). However, despide of these advancements, the automatic validation of both the correctness and performance of optimizations remains underexplored.
The primary objective of PCAOT is to automatically validate the correctness and performance of optimizations generated by Automatic Optimization Tools (AOTs). Our work aims to fill this gap by providing a comprehensive framework for evaluating and validating the outputs of AOTs, ensuring that the generated optimizations meet the desired standards of correctness. To the best of our knowledge, no other research has directly addressed this specific objective.
Last but not least, several benchmarks have been proposed to evaluate the capacity of AOTs on their abilities and limitations in optimizing code for maximum efficiency  (Bailey et al., 1991),  (Yuki and Pouchet, [n. d.]), (Black et al., 2022). We proposed a novel Parallel Computing Challenge Benchmark suite (PCB) V1.0 containing 28 use cases, each representing one of the six challenge patterns for evaluating AOT code optimization capabilities. To the best of our knowledge, no other benchmark have been designed to directly addressed this specific goal.
@@ 6. Conclusions
This paper addresses an existential question of the compiler community: ”Can AI-driven models revolutionize the way we approach code optimization?” To this end we evaluated five different Automatic Optimization Tools (AOT - three classical compilers and two LLMs), assessing their respective abilities and limitations in optimizing code for maximum efficiency. To support this evaluation, we introduced a novel mechanism called PCAOT, which enables the automatic validation of the correctness and performance of code generated by AOTs. While we have applied the mechanism to verify LLM-generated code, it also enables unsafe compiler optimization to be tested. Additionally, we have presented the Parallel Computing Challenge Benchmark suite (PCB) v1.0, comprising 28 parallel and non-parallel use cases for evaluating the capabilities of AOTs. To the best of our knowledge, this is the first such benchmark.
Our results show that LLMs only succeed in correctly optimizing very small programs of approximately 20 lines. Given that real-world software often contains thousands of lines of code, these models are not yet ready to serve as automatic optimizers.
In terms of performance, the chosen LLMs can successfully identify three out of the six different challenge patterns, including Array Reduction, Parallelizing Outermost Loops, and Loops Containing Function Calls, with a success rate higher than 30% in most cases. However, they completely fail on other patterns, including eliminating barrier synchronization using the NOWAIT clause, parallel regions enclosing multiple parallel loops, and avoiding load imbalance through dynamic scheduling.
We found substantial differences between three prompting strategies we employed for both the parallel and non-parallel scenarios, with ”Detailed Instruction Prompting” (DIP) being better among the three different approaches in most of the cases. Additionally, ”Chain-of-Thought” (CoT) prompting performed essentially similarly to DIP. We attribute this behavior to the high-level compiler knowledge strategies expressed in these two prompts.
In terms of speedup, the best compiler and the best LLM performed within 10%. CodeLlama, which is specifically optimized for code synthesis and infilling/completion tasks (Grattafiori et al., 2023), was expected to demonstrate strong performance in this context. However, ChatGPT contains a larger number of AI model parameters, which may make up for the fact that it is a general model.
Finally, addressing the main question asked in the introduction, we find that, with their current capabilities, the LLMs cannot yet suitably replace classical compiler optimization technology. Even though the potential exists, future LLMs will need significant improvements in terms of both performance and correctness. Future work will explore additional factors, such as the energy consumption required to achieve performance gains over conventional optimizing compilers. We also anticipate that our results will remain consistent across different platforms.
Acknowledgements.